OpenStudy (anonymous):

is this lagrange? i dont even know test tomorrow HALP!! If f(x)= sum ((2x)^(k+1))/(k+1) then what is f'(-1/3)? I have some ideas, but what am i doing wrong?

OpenStudy (anonymous):

ok so for f'(x) i got 2+4x+x^2+16x^3....2(2x)^k for general term

OpenStudy (anonymous):

so do you just plug in (-1/3) into that equation?

OpenStudy (anonymous):

is f defined as \[ f(x)=\sum _{k=0}^{\infty } \frac{2^{k+1} x^{k+1}}{k+1} \]

OpenStudy (anonymous):

\[ f'(x)=\sum _{k=0}^{\infty } 2^{k+1} x^k \]

OpenStudy (anonymous):

\[ f'(-1/3)=\sum _{k=0}^{\infty } 2^{k+1} \left(-\frac{1}{3}\right)^k =\frac{6}{5} \]